Status Report

NOAA GOES SATOPS Morning Report 17 Jan 2001

By SpaceRef Editor
January 17, 2001
Filed under ,

SATOPS MORNING REPORT – January 17, 2001

POES

NOAA-16 CPU-A – On January 17, at approximately 1800Z, CPU A on NOAA-16
started responding anomalously.  Initial indications indicated a possible
mixed (decryption/bypass) mode on the spacecraft, but subsequent analysis
indicated that a subset of telemetry coming from the processor was no longer
updating.  Engineers supported operations on-site overnight and will
discuss possible actions with NASA and the spacecraft contractor today. 
No products were impacted by this activity.

DMSP

OPS47 rev 30008 / POGO at 1309Z on January 16:  Cat 1 data loss
declared by user.  This is the last of a series of data losses related
to problems with AFSCN tracking station command connectivity.

Thule Tracking Station (POGO) comm problem summary:  GE Americom
was notified at 1920Z, January 15, of commanding problem with circuit from
SOCC to POGO.  Job repair ticket opened.  Loose cable on equipment
at Thule was the problem.  Link operational as of 1230Z on January
16.

GOES

GOES-8 momentum wheel 2 current – On January 17, at approximately 22:40Z
(during single chord operations), momentum wheel #2 on GOES-8 operated
at a high sustained current of ~1.0 amps (normal operating range ~450mA)
for about 2.5 hours.  Engineers were on-site and were prepared to
support operations for potential contingency actions, but the wheel returned
back to the normal operating range shortly after the return to dual chord
operations at 1:07Z.  No products were impacted by this activity.

At 1748Z and 1808Z on January 16:  GOES-10 image frame break due
to high bit error rate at CDA.

At 1019Z on January 16:  GOES-8 image frame break due to loss of
signal synch.

 

Questions or comments to [email protected]

SpaceRef staff editor.